MONT 104N – Modeling the Environment Study Questions on Frankenstein, by Mary Shelley August 15, 2018

0. Look at the title page of the novel. What is the reference to Paradise Lost? Who wrote that and why is it relevant? Who was Prometheus and why is he relevant?

1. The narrative structure of Frankenstein is quite complex in that the “main story” is both framed by a different story line told in letters from Robert Walton to his sister, and then divided between narrations delivered by Victor Frankenstein to Robert Walton and by Frankenstein’s creature to Frankenstein himself. How does this all fit together? Is it an effective way to tell the story or is it just confusing with so many different voices recounting parts of the action? Why do you suppose Mary Shelley chose to tell the story that way?

2. How are Walton and Frankenstein connected? How are they similar? How are they different? What do think Frankenstein represents for Walton? Why do you suppose Walton is so lonely when he has a ship full of other men as potential companions? How are his choices impacted by what he learns from Frankenstein?

3. What does Frankenstein say about his early enthusiasms for Cornelius Agrippa, Paracelsus, and Albertus Magnus? Who were these men? What did they do? Why were they exercise such a hold on the young Frankenstein’s imagination? Would we call someone like Paracelsus a scientist today? If not that, what? Did their work contribute in some way to modern science?

4. What do you think about Victor Frankenstein as a student and a scientist? What about his relationships with the two teachers Krempe and Waldman? What about his friendship with Henry Clerval? Do you admire his goals? Do you dislike them? Explain.

5. What is Frankenstein’s purpose in pursuing science? What is his relationship to the natural world? What does he study?

6. What is “galvanism” and who is it named after? This was a relatively recent scientific discovery at the time (1818) that Mary Shelley was writing. How does she use the real science of her time in the story?

7. How would you interpret Frankenstein’s initial response to the success of his experiment? (See pp. 45–46.) Why do we equate beauty and goodness, or ugliness and evil?

8. Why are there so many references to sickness and fever in Frankenstein? Trace these references throughout the novel. What broader theme might Shelley be expressing?

9. Who is Elizabeth Lavenza and how does Frankenstein feel about her? What does their relationship tell you about Frankenstein’s values and personality?

10. How does Frankenstein react to the realization that his creature was probably responsible for the death of his younger brother William and then for the death of Justine as well. Is he partly responsible for the creature’s actions?

11. How does the creature get to know the de Lacey family in the cottage? What turns the creature irrevocably toward evil? What makes their rejection so painful?

12. One would have to say that the creature exhibits a high degree of intelligence because he essentially learns to read on his own by poring over books of The Sorrows of Young Werther, Paradise Lost, and Plutarch’s Lives that he finds abandoned. He obviously has strong feelings and he thinks he would be capable of love. Why is he treated so hideously? What does that say about the rest of humanity?

13. How does the creature himself explain his evil behavior when he makes contact with Frankenstein? (See Chapter II in Volume II.) How does he compare himself with Adam and Satan in Paradise Lost?

14. How are the events in the novel different from those in film adaptations of the story you might have seen? For example, how is the possible “bride” for the creature treated in the novel? Is this a climax or turning point in the story? How does the creature get back at Frankenstein for the fact that he reneged on his promise?

15. Once Frankenstein’s family and all his hopes for a normal life have been destroyed, he and the creature embark on a mad chase across the ice sheets toward the North Pole. Why does Mary Shelley set this final phase of the novel there? What associations does that set up for readers?

16. Do you think readers are supposed to learn a lesson by reading this novel? What is it and do you feel as though you have learned something valuable by thinking about the themes Shelley brings into the story? If not, why is it not successful?